OpenStudy (yessenia27):

What is the solution to the equation 3x + 2(x − 9) = 8x + x − 14?

OpenStudy (acwbab):

Is this multiple choice or no?

OpenStudy (yessenia27):

Yes

OpenStudy (acwbab):

What are the choices?

OpenStudy (yessenia27):

-8 -1 1 8

OpenStudy (mellyg):

move all of the variables to one side, and then combine all the constants after you distribute the first part of the question.

OpenStudy (acwbab):

Doing what MellyG recommended, I'd say that the answer is B

alones (alones):

Okay simplify both sides of the equation 3x+2(x−9)=8x+x−14 5x−18=9x−14 than subtract 9x from both sides 5x−18−9x=9x−14−9x −4x−18=−14 and add 18 to both sides−4x−18+18=−14+18 −4x=4 than at last u divide both sides by -4. \(\frac{ -4x }{ -4 }\) = \(\frac{ 4 }{ -4 }\) x=−1